1/1

Erich Kästner,

Der September

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Author of this German poem is Erich Kästner (∗1899-02-23 – †1974-07-29).

Arg-568-1148 (2010-09-21 15:52:12)

http://www.poezio.net/version?poem-id=568&version-id=1148